OpenStudy (anonymous):

Find an equation of the line passing through the point(−2, 2) and through the point of intersection of the lines 2x − y + 1 = 0 and 3x + y − 16 = 0. Write your answer in the form y = mx + b.

OpenStudy (anonymous):

First, you need the point where the two lines cross. That's the point where both their x and y values are the same on both. Solve both lines as a system of equations, that'll get you the point of intersection. Next, use that point and (-2,2) to get the slope. And finally use the equation for a line. y-y1=m(x-x1) to get the equation of the line.

OpenStudy (anonymous):

Thank You so much! All the info is helpful!

OpenStudy (anonymous):

fyi-On the last equation substitute only x1 and y1 with a point that passes through the line, and substitute m, but not x and y. You will end up with an equation that has x and y on it. Solve for y to have in y=mx+b form

OpenStudy (anonymous):

Okay forsure!